# Permuton and local limits for the Luce model

### v1

@@ 1.2 Limits of random permutations: local and permuton convergence
The theory of permutons was first developed in [HKM+13]. Permuton convergence has been studied in various models, including Erdős–Szekeres permutations [Rom06], Mallows permutations [Sta09a], exponential families on permutations [Muk16], fixed pattern densities [KKRW20] and many others. For random pattern-avoiding permutations, the limiting permutons are often random [BBF+20, Bor21c]. In this paper, we further contribute by analyzing the case of Luce-distributed permutations.
@@ 1.3 An overview of our main results
  • Section˜3 is dedicated to the study of the local limit of Luce permutations. In ˜3.1, we establish the quenched Benjamini–Schramm convergence of Luce-distributed permutations, while in ˜3.2, we prove a stronger result: a central limit theorem for the number of consecutive occurrences of patterns with a Berry–Esseen type error estimate. Sections 3.2 and 3.3 provide applications of these results.
@@ 3.1 Statement of the main results
For the Luce model, the quenched Benjamini–Schramm convergence follows from the following result, whose proof is postponed to the end of the section.
@@
The assumption in (3.1) is a regularity assumption for the weights $\theta_{1},\dots,\theta_{n}$ that will be explicitly verified in certain examples in the next two subsections. For instance, if $\pi=21$, i.e. the case of descents, then (3.1) requires the existence of a finite limit for:
@@
$-p\mathopen{}\mathclose{{\left(\theta_{i+\pi^{-1}(1)},\dots,\theta_{i+\pi^{-1}(k)}}}\right)p\mathopen{}\mathclose{{\left(\theta_{i+h+\pi^{-1}(1)},\dots,\theta_{i+h+\pi^{-1}(k)}}}\right)\Bigg{)},$
@@
Before proving ˜3.1 and ˜3.2, we provide some applications, including explicit computations for the quantities appearing in these two theorems.
We consider the canonical example of Sukhatme weights $\theta_{i}=n-i+1$.
Lemma 3.3.
Fix $k\geq 1$ and $\pi\in\mathcal{S}_{k}$. Consider the case of Sukhatme weights $\theta_{i}=n-i+1$. Then
$\Lambda(\pi)=\lim_{n\to+\infty}\frac{1}{n}\sum_{i=0}^{n-k}p\mathopen{}\mathclose{{\left(\theta_{i+\pi^{-1}(1)},\dots,\theta_{i+\pi^{-1}(k)}}}\right)=\frac{1}{k!}.$
Proof.
Note that $n-i-k+1\leq\theta_{i+\pi^{-1}(\ell)}\leq n-i$ for all $0\leq i\leq n-k$ and $\ell\in[k]$. Moreover,
$(k-\ell)(n-i)+k(-k+1)\\ \leq\mathopen{}\mathclose{{\left(\theta_{i+1}+\dots+\theta_{i+k}}}\right)-(\theta_{i+\pi^{-1}(1)}+\dots+\theta_{i+\pi^{-1}(\ell)})\leq\\ (k-\ell)(n-i)-\ell(-k+1),$
for all $0\leq i\leq n-k$ and $\ell\in[k]$. As a consequence, by the definition of the function $p$ in (1.1), we get that
$\frac{(n-i-k+1)^{k}}{k!(n-i)^{k}+o((n-i)^{k})}\leq p\mathopen{}\mathclose{{\left(\theta_{i+\pi^{-1}(1)},\dots,\theta_{i+\pi^{-1}(k)}}}\right)\leq\frac{(n-i)^{k}}{k!(n-i)^{k}+o((n-i)^{k})}$ (3.2)
for all $0\leq i\leq n-\lfloor\sqrt{n}\rfloor$. Now, recalling that
$\Lambda(\pi):=\lim_{n\to+\infty}\frac{1}{n}\sum_{i=0}^{n-k}p\mathopen{}\mathclose{{\left(\theta_{i+\pi^{-1}(1)},\dots,\theta_{i+\pi^{-1}(k)}}}\right),$
we write
$\Lambda(\pi)=\lim_{n\to+\infty}\mathopen{}\mathclose{{\left(\frac{1}{n}\sum_{i=0}^{n-\lfloor\sqrt{n}\rfloor}p\mathopen{}\mathclose{{\left(\theta_{i+\pi^{-1}(1)},\dots,\theta_{i+\pi^{-1}(k)}}}\right)+O(1/\sqrt{n})}}\right).$
Finally, noting that from (3.2) we have that $p\mathopen{}\mathclose{{\left(\theta_{i+\pi^{-1}(1)},\dots,\theta_{i+\pi^{-1}(k)}}}\right)\to\frac{1}{k!}$ for all $0\leq i\leq n-\lfloor\sqrt{n}\rfloor$, we conclude from the Stolz–Cesàro theorem that $\Lambda(\pi)=\frac{1}{k!}$. ∎
Consequently, we deduce that if $\sigma_{n}$ is Luce-distributed with Sukhatme weights $\theta_{i}=n-i+1$, then it behaves locally as a uniform permutation. More precisely, we have the following result.
@@ 3.2 Application 1: permutations with Sukhatme weights are locally uniform
$\big{(}\operatorname{\widetilde{c\text{-}occ}}(\pi,\sigma_{n})\big{)}_{h\geq 1,\pi\in\mathcal{S}_{h}}\xrightarrow{P}\mathopen{}\mathclose{{\left(\frac{1}{h!}}}\right)_{h\geq 1,\pi\in\mathcal{S}_{h}}$
w.r.t. the product topology. In particular, $\sigma_{n}$ quenched Benjamini–Schramm converges.
Proof.
Since the limits are deterministic, pointwise convergence in probability is equivalent to convergence in probability in the product topology. Hence, it suffices to show that for any fixed $h\geq 1$ and $\pi\in\mathcal{S}_{h}$,
$\operatorname{\widetilde{c\text{-}occ}}(\pi,\sigma_{n})\xrightarrow{P}\frac{1}{h!}.$
This follows by combining ˜3.3 and ˜3.1. ∎
One obtains the following technical result using the same types of arguments used for ˜3.3.
Lemma 3.5.
Fix $k\geq 1$ and $\pi\in\mathcal{S}_{k}$. Consider the case of Sukhatme weights $\theta_{i}=n-i+1$. Set for all $k\geq 1$ and $1\leq h\leq k-1$,
$\zeta_{\pi}(h,k):=\mathbb{P}\mathopen{}\mathclose{{\left(E_{\pi^{-1}(1)}<\dots<E_{\pi^{-1}(k)},\,\,E_{h+\pi^{-1}(1)}<\dots<E_{h+\pi^{-1}(k)}}}\right),$ (3.3)
where $(E_{i})_{i}$ are i.i.d. exponential random variables of parameter 1, and
$\nu_{\infty}(\pi)=\frac{1}{k!}+\frac{1-2k}{(k!)^{2}}+2\sum_{h=1}^{k-1}\zeta_{\pi}(h,k).$ (3.4)
Then, as $n\to\infty$,
$(\nu_{n}(\pi))^{2}\sim n\cdot\nu_{\infty}(\pi).$
We obtain the following stronger result by combining ˜3.5 and ˜3.2.
Theorem 3.6.
Let $\sigma_{n}$ be a Luce-distributed permutation of size $n$ with Sukhatme weights $\theta_{i}=n-i+1$. Fix $k\geq 1$ and $\pi\in\mathcal{S}_{k}$. Then
@@
We conclude by showing the explicit values for $\nu_{\infty}(\pi)$ for all patterns of size 2 and 3 (with some matching numerical simulations shown in Figure˜2):
@@
Note that these are the same values one obtains for uniform permutations. ˜3.6 is reminiscent of results showing that local statistics – such as descents – have limiting distributions that do not change between uniform and non-uniform models. For example, Kammoun [Kam22] proves such universality for the Mallows models, and Kim and Li [KL20] show that descents are asymptotically normal when restricted to permutations in a fixed conjugacy class. See also [Bor21a, GK23, FK24] for other similar universality results.
@@ 3.3 Application 2: permutations with exponential Sukhatme weights are locally Luce-distributed
Here, we consider the example of exponential Sukhatme weights $\theta_{i}=\alpha^{n-i+1}$ for some $\alpha>0$.

### v2

@@ 1.2 Limits of random permutations: local and permuton convergence
The theory of permutons was first developed in [HKM+13]. Permuton convergence has been studied in various models, including Erdős–Szekeres permutations [Rom06], Mallows permutations [Sta09a], exponential families on permutations [Muk16], fixed pattern densities [KKRW20] and many others. For random pattern-avoiding permutations, the limiting permutons are often random [MP14, BBF+20, Bor21c]. In this paper, we further contribute by analyzing the case of Luce-distributed permutations.
@@ 1.3 An overview of our main results
  • Section˜3 is dedicated to the study of the local limit of Luce permutations. In ˜3.1, we establish the quenched Benjamini–Schramm convergence of Luce-distributed permutations, while in ˜3.2, we prove a stronger result: a central limit theorem for the number of consecutive occurrences of patterns with a Berry–Esseen type error estimate. ˜3.3 ensures that the above two results hold under the same minimal assumptions on the weights that we have for the permuton limit. Sections 3.2 and 3.3 provide applications of these results.
@@ 3.1 Statement of the main results
For the Luce model, the quenched Benjamini–Schramm convergence follows from the following result, whose proof is postponed to Section˜3.4.
@@
The assumption in (3.1) is a regularity assumption for the weights $\theta_{1},\dots,\theta_{n}$ that will be verified in ˜3.3 below under a minimal assumption. For instance, if $\pi=21$, i.e. the case of descents, then (3.1) requires the existence of a finite limit for:
@@
$-p\mathopen{}\mathclose{{\left(\theta_{i+\pi^{-1}(1)},\dots,\theta_{i+\pi^{-1}(k)}}}\right)p\mathopen{}\mathclose{{\left(\theta_{i+h+\pi^{-1}(1)},\dots,\theta_{i+h+\pi^{-1}(k)}}}\right)\Bigg),$ (3.2)
@@
It is natural to wonder when the Assumption (3.1) in ˜3.1 holds. The next theorem shows that such assumption holds [^4] See Section 3.3 for another setting where this assumption still holds. as soon as the assumtion of ˜2.1 is satisfied, i.e. the one in our permuton limit result. It also explicitly compute the asymptotics for $(\nu_{n}(\pi))^{2}$ from ˜3.2.
@@
Set $f_{n}(y):=\theta_{\lfloor yn\rfloor}$ for all $y\in[0,1]$ and assume that $f_{n}\to f$ almost everywhere for some positive, finite, measurable function $f$ on $[0,1]$. Then, for all $k\geq 1$ and every permutation $\pi\in\mathcal{S}_{k}$,
$\Lambda(\pi)=\lim_{n\to+\infty}\frac{1}{n}\sum_{i=0}^{n-k}p\mathopen{}\mathclose{{\left(\theta_{i+\pi^{-1}(1)},\dots,\theta_{i+\pi^{-1}(k)}}}\right)=\frac{1}{k!},$ (3.3)
and
$\lim_{n\to\infty}\frac{(\nu_{n}(\pi))^{2}}{n}=\nu_{\infty}(\pi),$ (3.4)
where
$\nu_{\infty}(\pi)=\frac{1}{k!}+\frac{1-2k}{(k!)^{2}}+2\sum_{h=1}^{k-1}\zeta_{\pi}(h,k),$ (3.5)
with
$\zeta_{\pi}(h,k):=\mathbb{P}\!\mathopen{}\mathclose{{\left(E_{\pi^{-1}(1)}<\dots<E_{\pi^{-1}(k)},\;\;E_{h+\pi^{-1}(1)}<\dots<E_{h+\pi^{-1}(k)}}}\right),\quad\text{$1\leq h\leq k-1$,}$ (3.6)
where $(E_{j})_{j}$ are i.i.d. exponential random variables with parameter 1.
Consequently, thanks to ˜3.1 and ˜3.2, under the assumption of ˜3.3 on the weights $\theta_{1},\dots,\theta_{n}$, if $\sigma_{n}\sim\operatorname{Luce}(\theta_{1},\dots,\theta_{n})$, then it behaves locally as a uniform permutation, in the sense of quenched Benjamini–Schramm convergence, Moreover,
@@
Note that these are the same values one obtains for uniform permutations (see Section˜3.2 for even more explicit computations). Equation˜3.7 is reminiscent of results showing that local statistics – such as descents – have limiting distributions that do not change between uniform and non-uniform models. For example, Kammoun [Kam22] proves such universality for the Mallows models, and Kim and Li [KL20] show that descents are asymptotically normal when restricted to permutations in a fixed conjugacy class. See also [Bor21a, GK23, FK24] for other similar universality results.
Before proving ˜3.1, ˜3.2, and ˜3.3, we provide some applications, including explicit computations for the quantities appearing in these three theorems.
We consider the canonical example of Sukhatme weights $\theta_{i}=n-i+1$. In this case, $f_{n}(x)\to f(x)=1-x$ for all $x\in[0,1]$. Since $f(x)>0$ for a.e. $x$, we can apply ˜3.3 and obtain the following result.
Proposition 3.4.
$\big(\operatorname{\widetilde{c\text{-}occ}}(\pi,\sigma_{n})\big)_{k\geq 1,\pi\in\mathcal{S}_{k}}\xrightarrow{P}\mathopen{}\mathclose{{\left(\frac{1}{k!}}}\right)_{k\geq 1,\pi\in\mathcal{S}_{k}}$
w.r.t. the product topology. In particular, $\sigma_{n}$ quenched Benjamini–Schramm converges. Moreover, for every fixed $k\geq 1$ and $\pi\in\mathcal{S}_{k}$,
where $\nu_{\infty}(\pi)$ is defined as in (3.5).
We show the explicit values for $\nu_{\infty}(\pi)$ for all patterns of size 2 and 3 (with some matching numerical simulations shown in Figure˜2):
@@ 3.3 Application 2: permutations with exponential Sukhatme weights are locally Luce-distributed
Here, we consider the example of exponential Sukhatme weights $\theta_{i}=\alpha^{n-i+1}$ for some $\alpha>0$. Note that the assumption of ˜3.3 is not satisfied in this setting, but we will show that we can still prove local convergence for this model and a central limit theorem for consecutive patterns.
@@ 3.4 Proofs of the main results
Proof of ˜3.3.
Fix $k\in\mathbb{N}$. For a permutation $\pi$ in $\mathcal{S}_{k}$ define
$\Phi_{\pi}(x_{1},\dots,x_{k})\coloneqq\prod_{j=1}^{k}\frac{x_{\pi(j)}}{x_{\pi(j)}+x_{\pi(j+1)}+\cdots+x_{\pi(k)}}\,,\qquad(x_{1},\dots,x_{k})\in(0,\infty)^{k}.$ (3.16)
We first prove (3.3). Since we are proving the result for any permutation $\pi$, recalling the definition of $p(\cdot)$ from (1.1), it is sufficient to prove that
$\lim_{n\to\infty}\frac{1}{n}\sum_{i=0}^{n-k}\Phi_{\pi}\!\bigl(\theta_{i+1},\dots,\theta_{i+k}\bigr)=\frac{1}{k!}.$ (3.17)
First, we observe some basic properties of $\Phi_{\pi}$. Each factor in $\Phi_{\pi}$ is at most $1$, hence $0\leq\Phi_{\pi}\leq 1$. Moreover $\Phi_{\pi}$ is $C^{1}$ on $(0,\infty)^{k}$ and simultaneous rescaling of all coordinates does not change $\Phi_{\pi}$. In particular, for any $c>0$,
$\Phi_{\pi}(c,\dots,c)=\prod_{j=1}^{k}\frac{1}{k-j+1}=\frac{1}{k!}.$
We now divide the rest of the proof into five main steps (the first four steps are dedicaded to the proof of (3.17), then the final fifth step is dedicated to the proof of (3.4) from the theorem statement).
Step 1: Restrict $f$ on a large set and gain regularity. Because $f$ is positive and finite a.e., there exists a sequence of sets
$E_{m}\coloneqq\Bigl\{y\in[0,1]:\tfrac{1}{m}\leq f(y)\leq m\Bigr\}$
with $|E_{m}|\uparrow 1$ as $m\to\infty$, where $|E_{m}|$ denotes the Lebesgue measure of $E_{m}$. Fix $m\in\mathbb{N}$ and $\eta>0$. By Lusin’s theorem, there exists a compact set $K^{\prime}\subset E_{m}$ with $|K^{\prime}|\geq|E_{m}|-\eta$ on which $f$ is continuous (hence uniformly continuous). By Egorov’s theorem, there exists a measurable set $K\subset K^{\prime}$ with $|K|\geq|K^{\prime}|-\eta$ such that $f_{n}\to f$ uniformly on $K$.
Summarizing, $K$ is measurable, $|K|\geq|E_{m}|-2\eta$, $f$ is uniformly continuous on $K$, and $f_{n}\to f$ uniformly on $K$. Also, on $K$ we have $\frac{1}{m}\leq f\leq m$.
Step 2: Good windows and their proportion. For $n\in\mathbb{N}$ set $y_{i}:=i/n$ for $i=0,\dots,n$. Call $i\in\{0,\dots,n-k\}$ good if
$y_{i},\ y_{i}+\tfrac{1}{n},\ \dots,\ y_{i}+\tfrac{k}{n}\ \in K.$
Let $G_{n}$ be the set of good indices, and $B_{n}$ its complement in $\{0,\dots,n-k\}$ ($B_{n}$ denotes the set of bad indices). We claim
$B_{n}\leq(k+1)\cdot\#\{0\leq j\leq n:\ y_{j}\notin K\}+k.$ (3.18)
Indeed, if $i$ is bad, then for some $r\in\{0,\dots,k\}$ we have $y_{i+r}\notin K$. Each point $y_{j}\notin K$ can serve as $y_{i+r}$ for at most $k+1$ different $i$. Finally, there are at most $k$ boundary indices to account for the range constraint $0\leq i\leq n-k$.
We would like now to claim that $\frac{1}{n}\#\{0\leq j\leq n:\ y_{j}\notin K\}$ converges to the Lebesgue measure of the complement of $K$ in $[0,1]$, but we need some care since we do not know that $\partial K$ has zero Lebesgue measure. Hence we need to take a small detour. By inner regularity, choose a finite union of closed intervals $F\subset K$ with $|K\setminus F|<\eta$. Then
$\{0\leq j\leq n:y_{j}\notin K\}\;\leq\;\#\{0\leq j\leq n:y_{j}\notin F\}.$
Since $F$ is a finite union of intervals, $\frac{1}{n}\#\{0\leq j\leq n:y_{j}\in F\}\to|F|$, hence $\limsup_{n\to\infty}\frac{1}{n}\#\{0\leq j\leq n:y_{j}\notin F\}\leq 1-|F|$. Combining with (3.18) gives
$\limsup_{n\to\infty}\frac{\#B_{n}}{n}\;\leq\;(k+1)\,(1-|F|).$
Because $|F|\geq|K|-\eta\geq|E_{m}|-3\eta$, we obtain
$\limsup_{n\to\infty}\frac{\#B_{n}}{n}\;\leq\;(k+1)\,\bigl(1-|E_{m}|+3\eta\bigr).$ (3.19)
Step 3: Uniform control of summands on good indices. Fix a good $i\in G_{n}$. For $r=1,\dots,k$ write
$x_{i,r}\ :=\ \theta_{i+r}\ =\ f_{n}\!\Bigl(y_{i}+\frac{r}{n}\Bigr).$
Fix now $\varepsilon^{\prime}>0$. Since the entire window $\bigl\{y_{i}+\frac{r}{n}:1\leq r\leq k\bigr\}\subset K$ by definition of good indices and $f_{n}\to f$ uniformly on $K$, we get that for $n$ large enough,
$\max_{1\leq r\leq k}\Bigl|x_{i,r}-f\!\Bigl(y_{i}+\frac{r}{n}\Bigr)\Bigr|\ \leq\ \varepsilon^{\prime}.$ (3.20)
Because $f$ is uniformly continuous on $K$ and $k/n\to 0$, we also get that for $n$ large enough,
$\max_{1\leq r\leq k}\Bigl|f\!\Bigl(y_{i}+\frac{r}{n}\Bigr)-f(y_{i})\Bigr|\ \leq\ \varepsilon^{\prime}.$ (3.21)
Combining (3.20) and (3.21), we obtain that for $n$ large enough,
$\max_{1\leq r\leq k}\bigl|x_{i,r}-f(y_{i})\bigr|\ \leq\ 2\varepsilon^{\prime}.$ (3.22)
Since $\frac{1}{m}\leq f\leq m$ on $K$, choosing $\varepsilon^{\prime}\leq\min\{\frac{1}{4m},\,\frac{m}{2}\}$ we obtain, for $n$ large enough,
$\frac{1}{2m}\ \leq\ x_{i,r}\ \leq\ 2m,\qquad\text{for all $r=1,\dots,k$}.$
Hence, for $n$ large enough, the vector $(x_{i,1},\dots,x_{i,k})$ lies in the compact rectangle $\mathcal{R}_{m}:=[\frac{1}{2m},\,2m]^{k}\subset(0,\infty)^{k}$. Because $\Phi_{\pi}$ is $C^{1}$ on $(0,\infty)^{k}$, it is Lipschitz on $\mathcal{R}_{m}$: There exists $L_{m}<\infty$ such that
$\bigl|\Phi_{\pi}(u)-\Phi_{\pi}(v)\bigr|\ \leq\ L_{m}\,\|u-v\|_{\infty},\qquad\text{ for all $u,v\in\mathcal{R}_{m}$}.$
Applying this with $u=(x_{i,1},\dots,x_{i,k})$ and $v=(f(y_{i}),\dots,f(y_{i}))$ gives (thanks to (3.22)) the following bound: for all $n$ large enough,
$\Bigl|\,\Phi_{\pi}\bigl(x_{i,1},\dots,x_{i,k}\bigr)-\Phi_{\pi}\bigl(f(y_{i}),\dots,f(y_{i})\bigr)\,\Bigr|\ \leq\ 2L_{m}\,\varepsilon^{\prime}.$ (3.23)
Since $\Phi_{\pi}(c,\dots,c)=1/k!$ for all $c>0$, (3.23) yields, for all good $i\in G_{n}$ and all $n$ large enough,
$\Bigl|\,\Phi_{\pi}\bigl(\theta_{i+1},\dots,\theta_{i+k}\bigr)-\tfrac{1}{k!}\,\Bigr|\ \leq\ 2L_{m}\,\varepsilon^{\prime}.$ (3.24)
Step 4: Average and limits. Using $0\leq\Phi_{\pi}\leq 1$ and decomposing the average over $G_{n}$ and $B_{n}$, we get
$\mathopen{}\mathclose{{\left|\frac{1}{n}\sum_{i=0}^{n-k}\Phi_{\pi}(\theta_{i+1},\dots,\theta_{i+k})-\frac{1}{k!}}}\right|\ \leq\ \frac{1}{n}\sum_{i\in G_{n}}\Bigl|\,\Phi_{\pi}(\theta_{i+1},\dots,\theta_{i+k})-\tfrac{1}{k!}\,\Bigr|+\frac{\#B_{n}}{n}.$
By (3.24), for $n$ large enough,
$\frac{1}{n}\sum_{i\in G_{n}}\Bigl|\,\Phi_{\pi}(\theta_{i+1},\dots,\theta_{i+k})-\tfrac{1}{k!}\,\Bigr|\ \leq\ 2L_{m}\,\varepsilon^{\prime}\,\frac{\#G_{n}}{n}\ \leq\ 2L_{m}\,\varepsilon^{\prime},$
since $G_{n}\subset\{0,\dots,n-k\}$ by definition. Taking $\varepsilon^{\prime}:=\min\{\frac{1}{4m},\,\frac{m}{2},\,\frac{\varepsilon}{4L_{m}}\}$ gives
$\limsup_{n\to\infty}\mathopen{}\mathclose{{\left|\frac{1}{n}\sum_{i=0}^{n-k}\Phi_{\pi}(\theta_{i+1},\dots,\theta_{i+k})-\frac{1}{k!}}}\right|\ \leq\ \frac{\varepsilon}{2}+\limsup_{n\to\infty}\frac{\#B_{n}}{n}.$
By (3.19), we get that
$\limsup_{n\to\infty}\mathopen{}\mathclose{{\left|\frac{1}{n}\sum_{i=0}^{n-k}\Phi_{\pi}(\theta_{i+1},\dots,\theta_{i+k})-\frac{1}{k!}}}\right|\ \leq\ \frac{\varepsilon}{2}+(k+1)\,(1-|E_{m}|+3\eta).$
Now, first let $\eta\downarrow 0$, then let $m\to\infty$ so that $|E_{m}|\uparrow 1$, and finally, let $\varepsilon\downarrow 0$. This concludes the proof of (3.17).
Step 5: The limiting variance. We finally prove the claim in (3.4). Set for all $h\in\{1,\ldots,k-1\}$ and $(x_{1},\dots,x_{h+k})\in(0,\infty)^{h+k}$,
$J^{h}_{\pi}(x_{1},\dots,x_{h+k})\coloneqq\mathbb{P}\mathopen{}\mathclose{{\left(E_{\pi(1)}<\dots<E_{\pi(k)},E_{h+\pi(1)}<\dots<E_{h+\pi(k)}}}\right),$
where $(E_{i})_{i}$ are independent exponential random variables of parameter $(x_{i})_{i}$. Recall also the function $\Phi_{\pi}$ introduced in (3.16).
Again, since we are proving the result for any permutation $\pi$, it is sufficient to prove that (recall the expression for $(\nu_{n}(\pi))^{2}$ from (3.2) in ˜3.2)
$\frac{1}{n}\sum_{i=0}^{n-k}\mathopen{}\mathclose{{\left(\Phi_{\pi}\!\bigl(\theta_{i+1},\dots,\theta_{i+k}\bigr)-\Phi_{\pi}\!\bigl(\theta_{i+1},\dots,\theta_{i+k}\bigr)^{2}}}\right)\\ +\frac{2}{n}\sum_{h=1}^{k-1}\sum_{i=0}^{n-k-h}\mathopen{}\mathclose{{\left(J^{h}_{\pi}\bigl(\theta_{i+1},\dots,\theta_{i+k}\bigr)-\Phi_{\pi}\!\bigl(\theta_{i+1},\dots,\theta_{i+k}\bigr)\Phi_{\pi}\!\bigl(\theta_{i+h+1},\dots,\theta_{i+h+k}\bigr)}}\right)\\ \longrightarrow\frac{1}{k!}+\frac{1-2k}{(k!)^{2}}+2\sum_{h=1}^{k-1}\zeta_{\pi^{-1}}(h,k),$
where $\zeta_{\pi}(h,k)$ was introduced in (3.6).
Note that for all $c>0$,
$\Phi_{\pi}(c,\ldots,c)=1/k!\qquad\text{and}\qquad J_{\pi}^{h}(c,\ldots,c)=\zeta_{\pi}(h,k).$
Moreover, for every $m\in\mathbb{N}$, there exists $L_{m}<\infty$ such that for all $h\in\{1,\ldots,k-1\}$,
$|\Phi_{\pi}(u)-\Phi_{\pi}(v)|\leq L_{m}\|u-v\|_{\infty}\quad\forall\,u,v\in[1/m,m]^{k},$
$|J^{h}_{\pi}(u)-J^{h}_{\pi}(v)|\leq L_{m}\|u-v\|_{\infty}\quad\;\forall\,u,v\in[1/m,m]^{h+k}.$ (3.25)
Indeed, the first inequality was already used in Step 3 of the proof and for the second inequality it is enought to notice the following facts: for $u=(u_{1},\dots,u_{h+k})\in[1/m,m]^{h+k}$,
$J^{h}_{\pi}(u)=\int_{A}\Big(\prod_{r=1}^{h+k}u_{r}\Big)\exp\!\Big(-\sum_{r=1}^{h+k}u_{r}t_{r}\Big)\,dt,$
where $A\subset(0,\infty)^{h+k}$ is a polyhedral cone. Differentiating the integrand, we obtain
$\frac{\partial}{\partial u_{r}}\Big[\prod_{s}u_{s}\,e^{-\sum_{s}u_{s}t_{s}}\Big]=\Big(\tfrac{1}{u_{r}}-t_{r}\Big)\prod_{s}u_{s}\,e^{-\sum_{s}u_{s}t_{s}}.$
For $u_{r}\in[1/m,m]$, this is bounded by $C_{m}(1+t_{r})e^{-1/m\sum_{s}t_{s}}$, which is integrable on $(0,\infty)^{h+k}$, uniformly in $u$. By dominated convergence, $J^{h}_{\pi}$ is $C^{1}$ on $1/m,m]^{h+k}$ and $\nabla J^{h}_{\pi}$ is bounded. Taking $L_{m}:=\sup_{h\in\{1,\dots,k-1\},\xi\in[1/m,m]^{h+k}}\|\nabla J^{h}_{\pi}(\xi)\|_{1}$ proves (3.25).
Since we have the same properties for $J^{h}_{\pi}$ as we had for $\Phi_{\pi}$, using exactly the same proof we used to prove in the previous steps that
$\lim_{n\to\infty}\frac{1}{n}\sum_{i=0}^{n-k}\Phi_{\pi}\!\bigl(\theta_{i+1},\dots,\theta_{i+k}\bigr)=\frac{1}{k!},$
we get that
$\frac{1}{n}\sum_{i=0}^{n-k}\mathopen{}\mathclose{{\left(\Phi_{\pi}\!\bigl(\theta_{i+1},\dots,\theta_{i+k}\bigr)-\Phi_{\pi}\!\bigl(\theta_{i+1},\dots,\theta_{i+k}\bigr)^{2}}}\right)\longrightarrow\frac{1}{k!}-\frac{1}{(k!)^{2}},$
and that
$\frac{2}{n}\sum_{h=1}^{k-1}\sum_{i=0}^{n-k-h}\mathopen{}\mathclose{{\left(J^{h}_{\pi}\bigl(\theta_{i+1},\dots,\theta_{i+k}\bigr)-\Phi_{\pi}\!\bigl(\theta_{i+1},\dots,\theta_{i+k}\bigr)\Phi_{\pi}\!\bigl(\theta_{i+h+1},\dots,\theta_{i+h+k}\bigr)}}\right)\\ \longrightarrow 2\sum_{h=1}^{k-1}\zeta_{\pi^{-1}}(h,k)-\frac{2k-2}{(k!)^{2}}.$
The sum of the right hand side of the last two equations is
$\frac{1}{k!}+\frac{1-2k}{(k!)^{2}}+2\sum_{h=1}^{k-1}\zeta_{\pi^{-1}}(h,k),$
as we wanted. This concludes the proof of (3.4), and thereby the entire theorem. ∎